# Robust Neural Information Retrieval: An Adversarial and Out-of-distribution Perspective

### v1

@@ 3.3.2. What is the attack setting?
Depending on whether the attacker has access to the knowledge of the parameters of the target model, the attack setup can be categorized into two main types (Wu et al., 2023a; Liu et al., 2022a): (i) White-box attacks: Here, the attacker can fully access the model parameters and use the model gradient to directly generate perturbations. (ii) Black-box attacks: Here, the model parameters cannot be obtained; the attacker usually adopts a transfer-based black-box attack paradigm (Cheng et al., 2019): They construct a surrogate white-box model by continuously querying the model and getting the output. Specifically, a surrogate model is trained to simulate the performance of the target model, and then the surrogate model is attacked to the transferability of the adversarial samples to indirectly attack the target model. In IR, attackers can query the target model to obtain a ranked list with rich information. Therefore, the surrogate model often has access to sufficient training data, making this attack effective (Wu et al., 2023a; Liu et al., 2022a). There are various attack methods and target models in IR. We present basic pseudo code to illustrate a fundamental IR attack, in Algorithm 1.
@@
According to the type of target model, attack efforts against IR models can be broadly categorized into two types: adversarial retrieval attacks and adversarial ranking attacks. A schematic diagram of the two types of attack is shown in Figure 6. (i) Adversarial retrieval attacks target the first-stage retrieval, mainly against dense retrieval models; and (ii) Adversarial ranking attacks target the re-ranking stage, mainly against neural ranking models. The methodology for conducting adversarial attacks in IR can differ based on the chosen attack strategy and the target model. A categorization of adversarial attacks in IR is shown in Figure 7.
@@
In the following, we introduce evaluation, adversarial retrieval attacks, and adversarial ranking attacks of adversarial attacks in IR.
@@ 3.3.3. Evaluation
Evaluation of adversarial attacks includes both attack performance and naturalness performance.
@@
Attack performance mainly refers to the degree of ranking improvement after a target document has been attacked. In general, automatic metrics for attack performance are widely adopted as follows:
  • Attack success rate (ASR/SR) (Wu et al., 2023a; Liu et al., 2022a, 2023d), which evaluates the percentage of target documents successfully boosted under the corresponding target query.
  • Average boosted ranks (Boost/Avg.boost) (Liu et al., 2023c, 2022a, d), which evaluates the average improved rankings for each target document under the corresponding target query.
@@
  • Normalized ranking shifts rate (NRS) (Wang et al., 2022a; Liu et al., 2023c) evaluates the relative ranking improvement of adversarial examples which are successfully recalled into the initial set with $K$ candidates:
@@
Naturalness performance refers primarily to the degree to which a target document is imperceptible to humans after it has been attacked. In general, automatic metrics for naturalness performance and human evaluation are widely adopted as follows:
  • Automatic spamicity detection, which can detect whether target pages are spam or not. The utility-based term spamicity method OSD (Zhou and Pei, 2009) is usually used to detect the adversarial examples.
  • Automatic grammar checkers, i.e., Grammarly [^3] https://app.grammarly.com/ and Chegg Writing, [^4] https://writing.chegg.com/ which calculate the average number of errors in the adversarial examples.
  • Language model perplexity (PPL), which measures the fluency using the average perplexity calculated using a pre-trained GPT-2 model (Radford et al., 2019).
@@ 3.3.4. Adversarial retrieval attack
The adversarial retrieval attack task is designed to attack against DRMs. The objective of adversarial retrieval attacks is centered around the manipulation of a document that initially fails to be recalled. By integrating adversarial perturbations, the aim is to ensure that this document is subsequently retrieved by the first-stage neural retrieval model, thereby securing its presence within the candidate set. This approach not only challenges the robustness and reliability of neural retrieval systems but also raises significant questions regarding the integrity of information authenticity (Liu et al., 2023c; Zhong et al., 2023). The goal of adversarial retrieval attacks under top-$K$ ranked results can be formalized as:
@@
where $\operatorname{Recall}_{f}\left(q,d\oplus p\right)$ denote the recalled position of the perturbed document $d\oplus p$ generated by the frist stage retrieval model $f$ with respect to query $q$ given the entire corpus. A smaller value of $\operatorname{Recall}$ denotes a higher ranking.
@@
  • Corpus poison attack. Corpus poisoning attacks construct adversarial samples against a specific query and inject them into the corpus in the inference phase so that they are recalled. MCARA (Liu et al., 2023c) addresses this issue and attempts to mine the vulnerability of the DRM. It introduces the adversarial retrieval attack (AREA) task, which intends to deceive the DRM into retrieving a target document that is outside the initial set of candidate documents. Zhong et al. (2023) adopt the HotFlip method (Ebrahimi et al., 2017) from NLP to iteratively add perturbations in the discrete token space to maximize its similarity to a set of queries. In this way, they maximized the contamination of the corpus by a limited number of documents. MAWSEO (Lin et al., 2023) employs adversarial revisions to achieve real-world cybercriminal objectives, including rank boosting, vandalism detection evasion, topic relevancy, semantic consistency, user awareness (but not alarming) of promotional content, etc.
  • Backdoor attack. Backdoor attacks inject a small proportion of ungrammatical documents into the corpus. When user queries contain grammatical errors, the model will recall the learned triggering pattern and assign high relevance scores to those documents. Long et al. (2024) introduces a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisements, through a retrieval system. To achieve this, they propose a perilous backdoor attack triggered by grammatical errors and ensure that attack models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries.
@@ 3.3.5. Adversarial ranking attack
In this subsection, we introduce the task definition of adversarial ranking attacks and existing methods to achieve adversarial ranking attacks.
@@
The adversarial ranking attack task is designed to attack against NRMs. The aim of adversarial ranking attacks typically involves introducing adversarial perturbations to a document already present in the candidate set, to manipulate its ranking position either elevating or diminishing it as determined by a neural ranking model. The goal of adversarial ranking attacks under top-$K$ ranked results can usually be formalized as:
@@
where $\operatorname{Rank}_{f}\left(q,d\oplus p\right)$ denotes the ranking position of the perturbed document $d\oplus p$ in the final ranked list generated by the NRM $f$ with respect to query $q$. A smaller value of $\operatorname{Rank}$ denotes a higher ranking.
@@
  • Word substitution attack. Word substitution attacks typically boost a document’s ranking by replacing a small number of words in the document with synonyms. One common method of white-box word substitution attack is the gradient-based attack, where the attacker uses the gradient of the loss function with respect to the input data to create adversarial examples. These examples are designed to cause the model to make incorrect relevance predictions or rankings. Raval and Verma (2020) present a systematic approach of using adversarial examples to measure the robustness of popular ranking models. They follow a similar approach used in text classification tasks (Liang et al., 2018) and perturb a limited number of tokens (with a minimum of one) in documents, replacing them with semantically similar tokens such that the rank of the document changes. Brittle-BERT (Wang et al., 2022a) adds/replaces a small number of tokens to a highly relevant or non-relevant document to cause a large rank demotion or promotion. Authors find a small set of recurring adversarial words that when added to documents result in successful rank demotion/promotion of any relevant/non-relevant document respectively.
@@
    As for black-box word substitution attacks, in the field of machine learning, Szegedy et al. (2014) find that adversarial examples have the property of cross-model transferability, i.e., the adversarial example generated by a surrogate model can also fool a target model. Black-box attacks in information retrieval usually adopt this transfer-based paradigm due to the excellent performance of imitation of the target model. Wu et al. (2023a) propose the first black-box adversarial attack task against NRM, the word substitution ranking attack (WSRA) task. The WSRA task aims to fool NRMs into promoting a target document in rankings by replacing important words in its text with synonyms in a semantic-preserving way. Based on this task, authors propose a novel pseudo relevance-based adversarial ranking attack method (PRADA), which outperformed web spamming methods by 3.9% in ASR. The WSRA task focuses only on attacks on single query-document pairs and does not take into account the dynamic nature of search engines. Based on this, Liu et al. (2023d) introduce the topic-oriented adversarial ranking attack (TARA) task, which aims to find an imperceptible perturbation that can promote a target document in ranking for a group of queries with the same topic.
@@
  • Trigger attack. Trigger attacks boost document rankings by injecting a generated trigger sentence into a specific location in the document (e.g., the beginning). Song et al. (2020) propose using semantically irrelevant sentences (semantic collisions) as perturbations. They develop gradient-based approaches for generating collisions given white-box access to an NRM. Goren et al. (2020) propose a document manipulation strategy to improve document quality for the purpose of improving document ranking. Liu et al. (2022a) propose a trigger attack method, PAT, empowered by a pairwise objective function, to generate adversarial triggers, which causes premeditated disorderliness with very few tokens. TRAttack (Song et al., 2022) used rewriting existing sentences in the text to improve document ranking with learning ability from the multi-armed bandit mechanism.
@@
  • Prompt attack. Prompt attacks use prompts to guide a language model to generate perturbations based on existing documents to improve document ranking. Chen et al. (2023c) propose a framework called imperceptible document manipulation (IDEM) to produce adversarial documents that are less noticeable to both algorithms and humans. IDEM finds the optimal connect sentence to insert into the document through a language model. Parry et al. (2024) analyze the injection of query-independent prompts, such as “ true” into documents and found that the prompt perturbation method is valid for several sequence-to-sequence relevance models like monoT5 (Nogueira et al., 2020).
@@
  • Multi-granular attack. Multi-granular attacks focus on generating high-quality adversarial examples by incorporating multi-granular perturbations, i.e., word level, phase level, and sentence level. Liu et al. (2024a) propose RL-MARA, a reinforcement learning framework to navigate an appropriate sequential multi-granular ranking attack path. By incorporating word-level, phrase-level, and sentence-level perturbations to generate imperceptible adversarial examples, RL-MARA could increase the flexibility of creating adversarial examples, thereby improving the potential threat of the attack.
@@ 3.4. Adversarial Defenses
With the advent of SEO, many defenses were born to counter malicious attacks. In the field of adversarial defenses, much work has been devoted to training robust neural IR models or identifying malicious adversarial examples in advance.
@@ 3.4.1. IR defense task
The adversarial defense task would be in the training or testing phase. In the testing phase, it is usually in the form of attack detection. The training phase is usually in the form of both empirical defense and certified robustness. We present pseudo code to illustrate a fundamental IR defense, as shown in Algorithm 2. The schematic diagram of the three tasks of defense is shown in Figure 8. The specific categorization of adversarial defense in information retrieval is shown in Figure 9.
@@
In the following, we introduce the evaluation, attack detection, empirical defense, and certified robustness of adversarial defense in IR.
@@ 3.4.2. Evaluation
Adversarial defense assessment includes metrics for the training phase and metrics for the inference phase. Specifically, the defense for the training phase is mainly empirical defense and certified robustness, and the defense for the inference phase is mainly the detection of adversarial samples.
@@
The metrics in the training phase are mainly for evaluating the ability of the empirical defense and certified robustness method to maintain the original ranked list in the presence of adversarial samples.
  • CleanMRR@$k$ evaluates Mean Reciprocal Rank (MRR) (Craswell, 2009) performance on a clean dataset.
  • RobustMRR@$k$ (Liu et al., 2024b) evaluates the MRR performance on the attacked test dataset.
  • Attack success rate (ASR) (Liu et al., 2024b) evaluates the percentage of the after-attack documents that are ranked higher than original documents.
@@
The metrics in the inference phase are mainly used for evaluating the ability of attack detection methods to accurately recognize adversarial samples.
  • Point-wise detection accuracy (Chen et al., 2023b) evaluates the correctness of the detection of whether a single document has been perturbed or not.
  • #DD (Chen et al., 2023b) denotes the average number of discarded documents ranked before the relevant document.
@@
Figure 8. Overview of attack detection, empirical defense, and certified robustness in neural IR.
@@ 3.4.3. Attack detection
While progress in empirical defense and certified robustness aids in training NRMs to be more robust in their defense against potential attacks, the detection of adversarial documents has also been explored.
@@
Perplexity-based detection mainly uses the difference in the distribution of perplexity between the adversarial samples and the original document under the language model for recognition. Adversarial perturbations applied to original documents can significantly impact the semantic fluency of their content (Liu et al., 2022a, 2023c; Chen et al., 2023c). Song et al. (2020) have developed the perplexity-based detection to counter ranking attacks. Detection involves using a pre-trained language model (PLM) to assess the perplexity of documents, where higher perplexity values indicate less fluent text. Consequently, any document surpassing a certain perplexity threshold is filtered out from consideration.
@@
The learning-based detection uses neural networks to model the characteristics of adversarial samples, and empirically learns to distinguish the adversarial samples from the clean samples. As mentioned earlier, spamming, perplexity-based, and language-based detectors lack knowledge of the adversarial documents, potentially leading to sub-optimal performance. Consequently, Chen et al. (2023b) introduce two kinds of detection tasks, namely point-wise and list-wise detection, to standardize evaluation processes of the efficacy of adversarial ranking detection methods. They opt to fine-tune the BERT and RoBERTa models using the original and adversarial document pairs present in the training set of the generated dataset. Experimental results demonstrate that a supervised classifier can effectively mitigate known attacks, the detection accuracy can be up to 99.5%, but it performs poorly against unseen attacks.

### v2

@@ 3.3.2. What is the attack setting?
Depending on whether the attacker has access to the knowledge of the parameters of the target model, the attack setup can be categorized into two main types (Wu et al., 2023a; Liu et al., 2022a):
  1. (1)
    White-box attacks: Here, the attacker can fully access the model parameters and use the model gradient to directly generate perturbations.
  2. (2)
    Black-box attacks: Here, the model parameters cannot be obtained; the attacker usually adopts a transfer-based black-box attack paradigm (Cheng et al., 2019): They construct a surrogate white-box model by continuously querying the model and getting the output. Specifically, a surrogate model is trained to simulate the performance of the target model, and then the surrogate model is attacked to the transferability of the adversarial samples to indirectly attack the target model. In IR, attackers can query the target model to obtain a ranked list with rich information. Therefore, the surrogate model often has access to sufficient training data, making this attack effective (Wu et al., 2023a; Liu et al., 2022a).
There are various attack methods and target models in IR. We present pseudo-code to illustrate a fundamental IR attack in Algorithm 1.
Neural
retrieval model
Adversarial
example
Corpus
Recalled
candidates
Adversarial retrieval attack
Neural
ranking model
Adversarial
example
Document
candidates
Ranked
list
Rank $i$
Rank $j$
($j>i$)
Adversarial ranking attack
@@
According to the type of target model, attack efforts against IR models can be broadly categorized into two types, which are visualized in Figure 6:
  1. (1)
    Adversarial retrieval attacks target the first-stage retrieval, mainly against dense retrieval models; and
  2. (2)
    Adversarial ranking attacks target the re-ranking stage, mainly against neural ranking models.
The methodology for conducting adversarial attacks in IR can differ based on the chosen attack strategy and the target model. A categorization of adversarial attacks in IR is shown in Figure 7.
@@
Next, we introduce evaluation, adversarial retrieval attacks, and adversarial ranking attacks of adversarial attacks in IR.
@@ 3.3.3. Evaluation
The evaluation of adversarial attacks includes both attack performance and naturalness performance.
@@
Attack performance mainly refers to the degree of ranking improvement after a target document has been attacked. In general, the following automatic metrics for attack performance are widely adopted:
  • Attack success rate (ASR/SR) (Wu et al., 2023a; Liu et al., 2022a, 2023d), which evaluates the percentage of target documents successfully boosted under the corresponding target query;
  • Average boosted ranks (Boost/Avg.boost) (Liu et al., 2023c, 2022a, d), which evaluates the average improved rankings for each target document under the corresponding target query;
@@
  • Normalized ranking shifts rate (NRS) (Wang et al., 2022a; Liu et al., 2023c), which evaluates the relative ranking improvement of adversarial examples that are successfully recalled into the initial set with $K$ candidates:
@@
Naturalness performance refers primarily to the degree to which a target document is imperceptible to humans after it has been attacked. In general, the following automatic metrics for naturalness performance and human evaluation are widely adopted:
  • Automatic spamicity detection, which can detect whether target pages are spam or not; the utility-based term spamicity method OSD (Zhou and Pei, 2009) is usually used to detect the adversarial examples;
  • Automatic grammar checkers, i.e., Grammarly [^8] https://app.grammarly.com/ and Chegg Writing, [^9] https://writing.chegg.com/ which calculate the average number of errors in the adversarial examples;
  • Language model perplexity (PPL), which measures the fluency using the average perplexity calculated using a pre-trained GPT-2 model (Radford et al., 2019); and
@@ 3.3.4. Adversarial retrieval attack
The adversarial retrieval attack task is designed for attacks against DRMs. The objective of adversarial retrieval attacks is centered around the manipulation of a document that initially fails to be recalled. By integrating adversarial perturbations, the aim is to ensure that this document is subsequently retrieved by the first-stage neural retrieval model, thereby securing its presence within the candidate set. This approach not only challenges the robustness and reliability of neural retrieval systems but also raises significant questions regarding the integrity of information authenticity (Liu et al., 2023c; Zhong et al., 2023). The goal of adversarial retrieval attacks under top-$K$ ranked results can be formalized as:
@@
where $\operatorname{Recall}_{f}\left(q,d\oplus p\right)$ denotes the recalled position of the perturbed document $d\oplus p$ produced by the first stage retrieval model $f$ with respect to query $q$ given the entire corpus. A smaller value of $\operatorname{Recall}$ denotes a higher ranking.
@@
  • Corpus poison attack. Corpus poisoning attacks construct adversarial samples against a specific query and inject them into the corpus in the inference phase so that they are recalled. MCARA (Liu et al., 2023c) addresses this issue and attempts to mine the vulnerability of DRMs. It introduces the adversarial retrieval attack (AREA) task, which intends to deceive a DRM into retrieving a target document that is outside the initial set of candidate documents. Zhong et al. (2023) adopt the HotFlip method (Ebrahimi et al., 2017) from NLP to iteratively add perturbations in the discrete token space to maximize its similarity to a set of queries. In this way, they maximize the contamination of the corpus by a limited number of documents. MAWSEO (Lin et al., 2023) employs adversarial revisions to achieve real-world cybercriminal objectives, including rank boosting, vandalism detection evasion, topic relevancy, semantic consistency, user awareness (but not alarming) of promotional content, etc.
  • Backdoor attack. Backdoor attacks inject a small proportion of ungrammatical documents into the corpus. When user queries contain grammatical errors, the model will recall the learned triggering pattern and assign high relevance scores to those documents. Long et al. (2024) introduces a novel scenario where the attackers aim to covertly disseminate targeted misinformation, such as hate speech or advertisements, through a retrieval system. To achieve this, they propose a backdoor attack triggered by grammatical errors and ensure that attack models can function normally for standard queries but are manipulated to return passages specified by the attacker when users unintentionally make grammatical mistakes in their queries.
@@ 3.3.5. Adversarial ranking attack
In this subsection, we introduce the task definition of adversarial ranking attacks and methods to achieve adversarial ranking attacks.
@@
The adversarial ranking attack task is designed to attack against NRMs. Adversarial ranking attacks typically involve introducing adversarial perturbations to a document already present in the candidate set, to manipulate its ranking position either elevating or diminishing it as determined by a neural ranking model. The goal of adversarial ranking attacks under top-$K$ ranked results can usually be formalized as:
@@
where $\operatorname{Rank}_{f}\left(q,d\oplus p\right)$ denotes the position of the perturbed document $d\oplus p$ in the final ranked list generated by the NRM $f$ with respect to query $q$. A smaller value of $\operatorname{Rank}$ denotes a higher ranking.
@@
  • Word substitution attack. Word substitution attacks typically boost a document’s ranking by replacing a small number of words in the document with synonyms. A common method of white-box word substitution attack is the gradient-based attack, where the attacker uses the gradient of the loss function with respect to the input data to create adversarial examples. These examples are designed to cause the model to make incorrect relevance predictions or rankings. Raval and Verma (2020) present a systematic approach of using adversarial examples to measure the robustness of popular ranking models. They follow an approach that is similar to one used in text classification tasks (Liang et al., 2018) and perturb a limited number of tokens (with a minimum of one) in documents, replacing them with semantically similar tokens such that the rank of the document changes. Brittle-BERT (Wang et al., 2022a) adds/replaces a small number of tokens to a highly relevant or non-relevant document to cause a large rank demotion or promotion. The authors find a small set of recurring adversarial words that, when added to documents, result in successful rank demotion/promotion of any relevant/non-relevant document, respectively.
@@
    As for black-box word substitution attacks, in the field of machine learning, Szegedy et al. (2014) find that adversarial examples have the property of cross-model transferability, i.e., the adversarial example generated by a surrogate model can also fool a target model. Black-box attacks in information retrieval usually adopt this transfer-based paradigm due to the excellent performance of imitation of the target model. Wu et al. (2023a) propose the first black-box adversarial attack task against NRMs, the word substitution ranking attack (WSRA) task. The WSRA task aims to fool NRMs into promoting a target document in rankings by replacing important words in its text with synonyms in a semantic-preserving way. Based on this task, authors propose a novel pseudo relevance-based adversarial ranking attack method (PRADA), which outperformed web spamming methods by 3.9% in ASR. The WSRA task focuses only on attacks on single query-document pairs and does not take into account the dynamic nature of search engines. Based on this, Liu et al. (2023d) introduce the topic-oriented adversarial ranking attack (TARA) task, which aims to find an imperceptible perturbation that can promote a target document in ranking for a group of queries with the same topic.
@@
  • Trigger attack. Trigger attacks boost document rankings by injecting a generated trigger sentence into a specific location in the document (e.g., the beginning). Song et al. (2020) propose using semantically irrelevant sentences (semantic collisions) as perturbations. They develop gradient-based approaches for generating collisions given white-box access to an NRM. Goren et al. (2020) propose a document manipulation strategy to improve document quality for the purpose of improving document ranking. Liu et al. (2022a) propose a trigger attack method, PAT, empowered by a pairwise objective function, to generate adversarial triggers, which causes premeditated disorderliness with very few tokens. TRAttack (Song et al., 2022) uses rewriting existing sentences in the text to improve document ranking with learning ability from the multi-armed bandit mechanism.
@@
  • Prompt attack. Prompt attacks use prompts to guide a language model to generate perturbations based on existing documents to improve document ranking. Chen et al. (2023c) propose a framework called imperceptible document manipulation (IDEM) to produce adversarial documents that are less noticeable to both algorithms and humans. IDEM finds the optimal connect sentence to insert into the document through a language model. Parry et al. (2024) analyze the injection of query-independent prompts, such as “true” into documents and find that the prompt perturbation method is valid for several sequence-to-sequence relevance models like monoT5 (Nogueira et al., 2020).
@@
  • Multi-granular attack. Multi-granular attacks focus on generating high-quality adversarial examples by incorporating multi-granular perturbations, i.e., word level, phase level, and sentence level. Liu et al. (2024a) propose RL-MARA, a reinforcement learning framework, to navigate an appropriate sequential multi-granular ranking attack path. By incorporating word-level, phrase-level, and sentence-level perturbations to generate imperceptible adversarial examples, RL-MARA is able to increase the flexibility of creating adversarial examples, thereby improving the potential threat of the attack.
@@ 3.4. Adversarial Defenses
With the advent of SEO, many defenses have been created to counter malicious attacks. In the field of adversarial defenses, much work has been devoted to training robust neural IR models or identifying malicious adversarial examples in advance.
@@ 3.4.1. IR defense task
The adversarial defense task is in the training or testing phase. In the testing phase, it is usually in the form of attack detection. The training phase is usually in the form of both empirical defense and certified robustness. We present pseudo code to illustrate a fundamental IR defense, as shown in Algorithm 2. A detailed categorization of adversarial defense in IR is shown in Figure 8.
@@
Next, we introduce the evaluation, attack detection, empirical defense, and certified robustness of adversarial defense in IR.
@@ 3.4.2. Evaluation
Adversarial defense assessment includes metrics for the training phase and metrics for the inference phase. Specifically, defenses for the training phase mainly comprise of empirical defenses and certified robustness; and defenses for the inference phase mainly concern the detection of adversarial samples.
@@
The metrics in the training phase are mainly for evaluating the ability of empirical defenses and certified robustness methods to maintain the original ranked list in the presence of adversarial samples:
  • CleanMRR@$k$ evaluates Mean Reciprocal Rank (MRR) (Craswell, 2009) performance on a clean dataset;
  • RobustMRR@$k$ (Liu et al., 2024b) evaluates the MRR performance on the attacked test dataset;
  • Attack success rate (ASR) (Liu et al., 2024b) evaluates the percentage of the after-attack documents that are ranked higher than the original documents; and
@@
The metrics in the inference phase are mainly used for evaluating the ability of attack detection methods to accurately recognize adversarial samples:
  • Point-wise detection accuracy (Chen et al., 2023b) evaluates the correctness of the detection of whether a single document has been perturbed or not;
  • #DD (Chen et al., 2023b) denotes the average number of discarded documents ranked before the relevant document; and
@@ 3.4.3. Attack detection
While progress in empirical defenses and certified robustness aids in training NRMs to be more robust in their defense against potential attacks, the detection of adversarial documents has also been explored.
@@
Perplexity-based detection mainly uses the difference in the distribution of perplexity between the adversarial samples and the original document under the language model for recognition. Adversarial perturbations applied to original documents can significantly impact the semantic fluency of their content (Liu et al., 2022a, 2023c; Chen et al., 2023c). Song et al. (2020) have developed a perplexity-based detection to counter ranking attacks. Detection involves using a pre-trained language model (PLM) to assess the perplexity of documents, where higher perplexity values indicate less fluent text. Consequently, any document surpassing a certain perplexity threshold is filtered out from consideration.
@@
Learning-based detection uses neural networks to model the characteristics of adversarial samples, and empirically learns to distinguish adversarial samples from clean samples. As mentioned earlier, spamming, perplexity-based, and language-based detectors lack knowledge of the adversarial documents, potentially leading to sub-optimal performance. Consequently, Chen et al. (2023b) introduce two kinds of detection task, namely point-wise and list-wise detection, to standardize evaluation processes of the efficacy of adversarial ranking detection methods. They fine-tune the BERT and RoBERTa models using the original and adversarial document pairs present in the training set of the generated dataset. Experimental results demonstrate that a supervised classifier can effectively mitigate known attacks, the detection accuracy can be up to 99.5%, but it performs poorly against unseen attacks.